OpenStudy (anonymous):

Find the slope of the line passing through the points (0,4) and (-6,7)

OpenStudy (katrinakaif):

What is the formula in finidng slope?

OpenStudy (anonymous):

I will have to look that up in my book

OpenStudy (katrinakaif):

y2-y1 ----- x2-x1

OpenStudy (katrinakaif):

Your given points are: (0,4) or (x1,y1) (-6,7) or (x2,y2) Can you plug it in the equation for me?

OpenStudy (anonymous):

I am looking through my book it is a pdf so still looking

OpenStudy (katrinakaif):

I listed the equation above: y2-y1 ---- x2-x1

OpenStudy (anonymous):

oh y2-y1 ---- x2-x1

OpenStudy (katrinakaif):

Yes, plug it in

OpenStudy (anonymous):

what do you mean plug it in, I am really terrible at math, so not understanding this

OpenStudy (katrinakaif):

Your given points are: (0,4) or (x1,y1) (-6,7) or (x2,y2) Can you plug it in the equation for me?

OpenStudy (anonymous):

well the numbers are (0, -4) and (-6,7) is that what you mean

OpenStudy (katrinakaif):

y2-y1 7 - 4 ----- = ------- x2-x1 -6 + 0

OpenStudy (anonymous):

Okay I see how you set it up so where do I go from here?

OpenStudy (katrinakaif):

Simplify this: 7 - 4 ----- -6 + 0 3 = ---- -6

OpenStudy (anonymous):

okay so do I do anything else after that?

OpenStudy (katrinakaif):

Do you think 3/-6 can simplify more?

OpenStudy (anonymous):

well 3 does go into 6

OpenStudy (katrinakaif):

And then?

OpenStudy (anonymous):

1 -- 3

OpenStudy (katrinakaif):

Don't forget the negative! :-)

OpenStudy (anonymous):

okay so 1 --- -3

hero (hero):

@KatrinaKaif , I recommend using the equation editor.

OpenStudy (katrinakaif):

Normally, the negative is in the numerator. So -1/3

OpenStudy (anonymous):

Equation editor is that a website?

OpenStudy (katrinakaif):

@Hero will explain.

hero (hero):

Nope, it is a button that you click that enables you to create things like this: \[m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

wow, very cool. Is that on this page to use?

hero (hero):

There is a button named "Equation" underneath the typing box. Click that and see what happens. If you get the far, copy and paste this into it: m = frac{y_2 - y_1}{x_2 - x_1} Then click insert and post.

OpenStudy (anonymous):

\[I got \it \] haha and I love it thanks

OpenStudy (anonymous):

could not coy and paste though kept clicking off

hero (hero):

You have to copy it first, then open the equation editor, then paste. Try it :D

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay when I click insert after copy and paste it goes away

hero (hero):

Try it again, except this time, press the space bar once, then wait a few seconds. Make sure you click insert, then click post.

OpenStudy (anonymous):

\[m = \frac{y_2 - y_1}{x_2 - x_1} \]

hero (hero):

:D

OpenStudy (anonymous):

:)

hero (hero):

Basically, what I have given you is the format for posting fractions. So now you know how to post those.

OpenStudy (anonymous):

Thank you, you are my hero :)

hero (hero):

:D I get that a lot, thanks :D

OpenStudy (anonymous):

Seriously though, I really appreciate you taking the time to show my how to do that.

hero (hero):

No problem :D

OpenStudy (anonymous):

I have another problem I was going to try to type in but I don't see the equation editor at the "ask a ?" box

hero (hero):

You would have to put it in what's called "Latex" form in order for it to post in that manner.

OpenStudy (anonymous):

oh maybe i can type it in here and copy and paste unless you can help me with the problem?

hero (hero):

See your fan testimonial for an explanation.

OpenStudy (anonymous):

Okay thank you again hero :)

OpenStudy (anonymous):

\[y = -2x ^{2 + 12x - 13} \]